# Through the Looking Glass: Mirror Schrödinger Bridges

### v1

@@ 6 Conclusion
Figure 5: The effect of using small $\sigma$ when pushing samples forward. From left to right: initial samples, intermediate times, and samples at terminal time.
@@
By studying an overlooked version of the Schrödinger bridge problem, which we coin the mirror Schrödinger bridge, we present an algorithm to sample with control over the in-distribution variation of new data points. Our method is flexible and requires fewer training iterations than existing alternatives (De Bortoli et al., 2021; Shi et al., 2023) designed for the general Schrödinger bridge problem. From a theoretical perspective, our method presents advantages over mirror interpolants (Albergo et al., 2023), specifically by obtaining kinetic optimality. While one might consider optimizing fixed mirror interpolants, the resulting $\min$-$\max$ optimization problem is intractable (Shaul et al., 2023). By contrast, our method is numerically tractable, is well-principled, and cuts down training in applications where control over in-distribution variation is desired. On the application front, we demonstrate that our method is a flexible tool to obtain new data points from empirical distributions in a variety of domains, including $2$-dimensional measures and image datasets.
Acknowledgments
@@
The authors would like to thank Lingxiao Li, Artem Lukoianov and Christopher Scarvelis for discussion and feedback. We are also grateful to Suvrit Sra and Francisco Vargas for thoughtful insights on related work. We thank Ahmed Mahmoud for proofreading. Leticia Mattos Da Silva acknowledges the generous support of a MathWorks Engineering Fellowship. Silvia Sellán is supported by an MIT Postdoctoral Fellowship for Engineering Excellence. The MIT Geometric Data Processing Group acknowledges the generous support of Army Research Office grants W911NF2010168 and W911NF2110293, of National Science Foundation grant IIS-2335492, from the CSAIL Future of Data program, from the MIT–IBM Watson AI Laboratory, from the Toyota–CSAIL Joint Research Center, and from the Wistron Corporation.
Appendix A Proof of Proposition 3
The first claimed expression is the very definition of ${\mathbb{Q}}_{\rm{SB}}$. As for the second claimed expression, let ${\mathbb{Q}}^{k}$ be the sequence of IPFP iterates. Note that by Lemma 2, we have
${\mathbb{Q}}^{2k+1}=\operatorname*{arg\,min}_{{\mathbb{P}}\in{\mathbb{D}}(\cdot,\pi_{1})}D_{\mathrm{KL}}\left({\mathbb{Q}}^{2k}\;\|\;{\mathbb{Q}}\right),\;{\mathbb{Q}}^{2k+2}=\operatorname*{arg\,min}_{{\mathbb{P}}\in{\mathbb{D}}(\pi_{0},\cdot)}D_{\mathrm{KL}}\big{(}{\mathbb{Q}}^{2k+1}\;\|\;{{\mathbb{Q}}}\big{)}.$
Since ${\mathbb{Q}}_{\rm{SB}}$ belongs to both projection sets, the Pythagorean theorem for reverse $D_{\mathrm{KL}}$ projections (Brekelmans & Neklyudov, 2023, Theorem 3.4) (see also (Csiszár & Matus, 2003, Theorem 5)) yields that for each $k$ we have
@@ Appendix A Proof of Proposition 3
$D_{\mathrm{KL}}\left({\mathbb{Q}}^{0}\;\|\;{\mathbb{Q}}_{\rm{SB}}\right)=\sum_{i=1}^{k}D_{\mathrm{KL}}\left({\mathbb{Q}}^{i-1}\;\|\;{\mathbb{Q}}^{i}\right)+D_{\mathrm{KL}}\left({\mathbb{Q}}^{k}\;\|\;{\mathbb{Q}}_{\rm{SB}}\right)$ (13)
Now the sequence $D_{\mathrm{KL}}({\mathbb{Q}}^{k}\;\|\;{\mathbb{Q}}_{\rm{SB}})$ converges to zero because the sequence $D_{\mathrm{KL}}({\mathbb{Q}}_{\rm{SB}}\;\|\;{\mathbb{Q}}^{k})$ converges to zero (see, e.g., (Weis, 2014, Theorem 3.21.4)). Thus, taking the limit as $k\to\infty$, we deduce that
@@
$D_{\mathrm{KL}}\left({\mathbb{Q}}^{0}\;\|\;{\mathbb{Q}}_{\rm{SB}}\right)=\sum_{i=1}^{\infty}D_{\mathrm{KL}}\left({\mathbb{Q}}^{i-1}\;\|\;{\mathbb{Q}}^{i}\right)$
Now, write ${\mathbb{Q}}^{\star}=\operatorname*{arg\,min}_{{\mathbb{Q}}\in{\mathbb{D}}(\pi_{0},\pi_{1})}D_{\mathrm{KL}}\left({\mathbb{Q}}^{0}\;\|\;{\mathbb{Q}}\right)$. A similar argument shows that
$D_{\mathrm{KL}}\left({\mathbb{Q}}^{0}\;\|\;{\mathbb{Q}}^{\star}\right)$ $=\sum_{i=1}^{\infty}D_{\mathrm{KL}}\left({\mathbb{Q}}^{i-1}\;\|\;{\mathbb{Q}}^{i}\right)+\lim_{k\to\infty}D_{\mathrm{KL}}\left({\mathbb{Q}}^{k}\;\|\;{\mathbb{Q}}^{\star}\right)$
$=D_{\mathrm{KL}}\left({\mathbb{Q}}^{0}\;\|\;{\mathbb{Q}}_{\rm{SB}}\right)+\lim_{k\to\infty}D_{\mathrm{KL}}\left({\mathbb{Q}}^{k}\;\|\;{\mathbb{Q}}^{\star}\right)\geq D_{\mathrm{KL}}\left({\mathbb{Q}}^{0}\;\|\;{\mathbb{Q}}_{\rm{SB}}\right),$
where the last inequality above follows from the nonnegativity of the KL divergence. It follows that $D_{\mathrm{KL}}\left({\mathbb{Q}}^{0}\;\|\;{\mathbb{Q}}_{\rm{SB}}\right)$ also achieves the desired minimum $D_{\mathrm{KL}}$, i.e., we have $D_{\mathrm{KL}}\left({\mathbb{Q}}^{0}\;\|\;{\mathbb{Q}}_{\rm{SB}}\right)=D_{\mathrm{KL}}\left({\mathbb{Q}}^{0}\;\|\;{\mathbb{Q}}^{\star}\right)$. Finally, we must rule out the possibility that this minimizer is not unique. To do this, observe that, by the squeeze theorem, we must have
$\lim_{k\to\infty}D_{\mathrm{KL}}\left({\mathbb{Q}}^{k}\;\|\;{\mathbb{Q}}^{\star}\right)=0.$
We can now apply Pinsker’s Inequality, which tells us that the KL divergence $D_{\mathrm{KL}}$ is at least a constant multiple of the square of the metric $\delta_{\rm{TV}}$ induced by total variation. More precisely,we have that $D_{\mathrm{KL}}\left({\mathbb{Q}}^{k}\;\|\;{\mathbb{Q}}^{\star}\right)\geq 2\delta_{\rm{TV}}\left({\mathbb{Q}}^{k},{\mathbb{Q}}^{\star}\right)^{2}$. We deduce that
$\lim_{k\to\infty}\delta_{\rm{TV}}\left({\mathbb{Q}}^{k},{\mathbb{Q}}^{\star}\right)=0,$
which implies that ${\mathbb{Q}}^{k}$ converges to ${\mathbb{Q}}^{\star}$ in total variation. We conclude that ${\mathbb{Q}}^{\star}={\mathbb{Q}}_{\rm{SB}}$. ∎
Appendix B Analytical Solution for Gaussian Experiment
Proposition 4.

### v2

@@ 6 Conclusion
By studying an overlooked version of the Schrödinger bridge problem, which we coin the mirror bridge, we present an algorithm to sample with control over the in-distribution variation of new data points. Our method is flexible and requires fewer training iterations than existing alternatives (De Bortoli et al., 2021; Shi et al., 2023) designed for the general Schrödinger bridge problem. As a technique for estimating Schrödinger bridges in the mirror case, our method presents an advantage over mirror interpolants (Albergo et al., 2023), since optimizing the relative entropy of such an interpolant results in a $\min$-$\max$ optimization problem that is typically intractable (Shaul et al., 2023). Our method is numerically tractable and well-principled, and it cuts down training in applications where control over in-distribution variation is desired. On the application front, we demonstrate that our method is a flexible tool to obtain new data points from empirical distributions in a variety of domains, including $2$-dimensional measures and image datasets. In future work, we hope to study of a potential $\sigma$ threshold for a sample to change class when resampled or to make “class” a neural network input, similar to text prompting in image generation.
Acknowledgments and Disclosure of Funding
@@
The authors would like to thank Lingxiao Li, Artem Lukoianov and Christopher Scarvelis for discussion and feedback. We are also grateful to Suvrit Sra for thoughtful insights on related work. We thank Ahmed Mahmoud for proofreading. Leticia Mattos Da Silva acknowledges the generous support of a MathWorks Engineering Fellowship. The MIT Geometric Data Processing Group acknowledges the generous support of Army Research Office grants W911NF2010168 and W911NF2110293, of National Science Foundation grant IIS2335492, from the CSAIL Future of Data program, from the MIT–IBM Watson AI Laboratory, from the Wistron Corporation, and from the Toyota–CSAIL Joint Research Center.
Appendix A Proof of Proposition 4.4
As before, we are interested in computing the associated time-symmetric drift, rather than the path measure itself. To this end, let $\pi$ be a probability distribution on $\mathbb{R}^{n}$, and let ${\mathbb{P}}\in{\mathbb{S}}$ and ${\mathbb{P}}^{\dagger}\in{\mathbb{D}}(\pi,\cdot)$ be path measures corresponding to diffusion processes. Suppose we seek to minimize $D_{\mathrm{KL}}({\mathbb{P}}\;\|\;{\mathbb{P}}^{\dagger})$ over all ${\mathbb{P}}\in{\mathbb{S}}$. Let $a_{t}^{\pm}=\frac{1}{2}(f_{t}^{{\mathbb{P}}^{\dagger}}\pm b_{t}^{{\mathbb{P}}^{\dagger}})$, and recall from (10) that we have the following expression:
$D_{\mathrm{KL}}({\mathbb{P}}\;\|\;{\mathbb{P}}^{\dagger})$ $=C+\frac{1}{4\sigma^{2}}\int_{0}^{1}{\mathbb{E}_{{\mathbb{P}}}}\bigg{[}\big{(}v_{t}^{{\mathbb{P}}}({\mathbf{X}}_{t})-f_{t}^{{\mathbb{P}}^{\dagger}}({\mathbf{X}}_{t})\big{)}^{2}+\big{(}v_{t}^{{\mathbb{P}}}({\mathbf{X}}_{t})-b_{t}^{{\mathbb{P}}^{\dagger}}({\mathbf{X}}_{t})\big{)}^{2}\bigg{]}{\rm{d}}t$
$=C+\frac{1}{2\sigma^{2}}\int_{0}^{1}{\mathbb{E}_{{\mathbb{P}}}}\bigg{[}\big{(}v_{t}^{{\mathbb{P}}}({\mathbf{X}}_{t})-a_{t}^{+}({\mathbf{X}}_{t})\big{)}^{2}+a_{t}^{-}({\mathbf{X}}_{t})^{2}\bigg{]}{\rm{d}}t,$
where $C$ is a constant. We now apply calculus of variations to compute the desired series expansion for the drift of the minimizer of $D_{\mathrm{KL}}({\mathbb{P}}\;\|\;{\mathbb{P}}^{\dagger})$ over ${\mathbb{P}}\in{\mathbb{S}}$. To take the first variation of this $D_{\mathrm{KL}}$ divergence with respect to $v_{t}^{{\mathbb{P}}}$, we apply the product rule, first differentiating the argument of the expectation, and then differentiating the expectation operator itself. The calculation of the first variation of this sort of $D_{\mathrm{KL}}$ divergence has been performed before in the literature; see, e.g., [Nüsken and Richter, 2021, section A.2, proof of Proposition 4.3] for the full details. We reproduce the calculation here for the sake of completeness.
Let $\epsilon>0$, and let $h$ be a test function. Let ${\mathbb{P}}_{\epsilon}$ denote the path measure corresponding to the perturbed drift function $v_{t}^{{\mathbb{P}}_{\epsilon}}\coloneqq v_{t}^{{\mathbb{P}}}+\epsilon h$. Then using the Radon-Nikodym derivative ${\rm{d}}{\mathbb{P}}_{\epsilon}/{\rm{d}}{\mathbb{P}}$ to rewrite the expectation in ${\mathbb{P}}_{\epsilon}$ in terms of an expectation in ${\mathbb{P}}$, we find that
@@
$\left.\frac{\partial}{\partial\epsilon}\mathbb{E}_{{\mathbb{P}}_{\epsilon}}\left[{\int_{0}^{1}}\big{(}v_{t}^{{\mathbb{P}}_{\epsilon}}({\mathbf{X}}_{t})-a_{t}^{+}({\mathbf{X}}_{t})\big{)}^{2}+a_{t}^{-}({\mathbf{X}}_{t})^{2}\mathrm{d}t\right]\right|_{\epsilon=0}=$ (15)
@@
$\quad\quad\quad\quad\left.\frac{\partial}{\partial\epsilon}\mathbb{E}_{{\mathbb{P}}}\left[\frac{{\rm{d}}{\mathbb{P}}_{\epsilon}}{{\rm{d}}{\mathbb{P}}}{\int_{0}^{1}}\big{(}v_{t}^{{\mathbb{P}}_{\epsilon}}({\mathbf{X}}_{t})-a_{t}^{+}({\mathbf{X}}_{t})\big{)}^{2}+a_{t}^{-}({\mathbf{X}}_{t})^{2}\mathrm{d}t\right]\right|_{\epsilon=0}=$
$\quad\quad\quad\quad\quad\quad\quad\quad\mathbb{E}_{{\mathbb{P}}}\Bigg{[}{\left.\frac{{\rm{d}}{\mathbb{P}}_{\epsilon}}{{\rm{d}}{\mathbb{P}}}\right|_{\epsilon=0}}\Bigg{(}{\int_{0}^{1}}2h({\mathbf{X}}_{t})\cdot\big{(}v_{t}^{{\mathbb{P}}}({\mathbf{X}}_{t})-a_{t}^{+}({\mathbf{X}}_{t})\big{)}\mathrm{d}t$
$\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad+{\int_{0}^{1}}\bigg{(}\big{(}v_{t}^{{\mathbb{P}}}({\mathbf{X}}_{t})-a_{t}^{+}({\mathbf{X}}_{t})\big{)}^{2}+a_{t}^{-}({\mathbf{X}}_{t})^{2}\bigg{)}\mathrm{d}t\cdot\left.\bigg{(}\frac{\partial}{\partial\epsilon}\log\frac{{\rm{d}}{\mathbb{P}}_{\epsilon}}{{\rm{d}}{\mathbb{P}}}\bigg{)}\right|_{\epsilon=0}\Bigg{)}\Bigg{]}.$
where the second step above follows from the Radon-Nikodym Theorem, and where the third step above follows from the product rule and the definition of the logarithmic derivative.
To simplify the right-hand side of (15), we first note that $({\rm{d}}{\mathbb{P}}_{\epsilon}/{\rm{d}}{\mathbb{P}})\mid_{\epsilon=0}=1$. Secondly, we can apply Girsanov’s Theorem to obtain a formula for ${\rm{d}}{\mathbb{P}}_{\epsilon}/{\rm{d}}{\mathbb{P}}$ as follows:
$\frac{{\rm{d}}{\mathbb{P}}_{\epsilon}}{{\rm{d}}{\mathbb{P}}}=\exp\left(-\epsilon\int_{0}^{1}h({\mathbf{X}}_{t}){\rm{d}}W_{t}-\frac{\epsilon^{2}}{2}\int_{0}^{1}|h({\mathbf{X}}_{t})|^{2}{\rm{d}}W_{t}\right).$ (16)
Differentiating the logarithm of (16) with respect to $\epsilon$ and setting $\epsilon=0$ then yields
$\bigg{(}\left.\frac{\partial}{\partial\epsilon}\log\frac{{\rm{d}}{\mathbb{P}}_{\epsilon}}{{\rm{d}}{\mathbb{P}}}\bigg{)}\right|_{\epsilon=0}=-\int_{0}^{1}h({\mathbf{X}}_{t}){\rm{d}}W_{t}.$ (17)
Let $T$ denote the linear map defined on test functions by $h\mapsto-\int_{0}^{1}h({\mathbf{X}}_{t}){\rm{d}}W_{t}$. Then combining the right-hand side of (15) with (17), we obtain the following expression:
$\int_{0}^{1}\mathbb{E}_{{\mathbb{P}}}\left[2h\cdot\big{(}v_{t}^{{\mathbb{P}}}({\mathbf{X}}_{t})-a_{t}^{+}({\mathbf{X}}_{t})\big{)}+\bigg{(}\big{(}v_{t}^{{\mathbb{P}}}({\mathbf{X}}_{t})-a_{t}^{+}({\mathbf{X}}_{t})\big{)}^{2}+a_{t}^{-}({\mathbf{X}}_{t})^{2}\bigg{)}\cdot T(h)\right]{\rm{d}}t.$ (18)
For the quantity in (18) to be zero for every possible choice of the test function $h$, we must have that the linear map
$h\mapsto 2h\cdot\big{(}v_{t}^{{\mathbb{P}}}({\mathbf{X}}_{t})-a_{t}^{+}({\mathbf{X}}_{t})\big{)}+\bigg{(}\big{(}v_{t}^{{\mathbb{P}}}({\mathbf{X}}_{t})-a_{t}^{+}({\mathbf{X}}_{t})\big{)}^{2}+a_{t}^{-}({\mathbf{X}}_{t})^{2}\bigg{)}\cdot T(h)$ (19)
is identically zero. Setting this map equal to zero and solving for $v_{t}^{{\mathbb{P}}}$ yields the desired first-order approximation $v_{t}^{{\mathbb{P}}}=a_{t}^{+}+O(||a_{t}^{-}||^{2})$. ∎
Appendix B Heuristic Bound on Our Estimation
To understand whether our AMP yields a viable estimate of the MSB, it would be useful to be able to quantify the difference between the limit of our AMP and the MSB. In this section, we provide conditions on the marginal distribution $\pi$ and on the reference path measure ${\mathbb{P}}^{0}$ under which we can explicitly bound the difference. The conditions we include here are modeled after the work of Vargas et al. [2021], who also develop schemes for approximating Schrödinger bridges; see, e.g., Conjecture 1 in their Appendix F.
To do this, we first introduce the following terminology: given a path measure ${\mathbb{P}}$, define the following three projections:
$\Pi({\mathbb{P}})=\operatorname*{arg\,min}_{{\mathbb{Q}}\in{\mathbb{D}}(\pi,\cdot)}D_{\mathrm{KL}}\left({\mathbb{Q}}\;\|\;{\mathbb{P}}\right)$ (direct)
$\Phi({\mathbb{P}})=\operatorname*{arg\,min}_{{\mathbb{Q}}\in{\mathbb{S}}}D_{\mathrm{KL}}\left({\mathbb{Q}}\;\|\;{\mathbb{P}}\right)$ (direct)
$\Psi({\mathbb{P}})=\operatorname*{arg\,min}_{{\mathbb{Q}}\in{\mathbb{S}}}D_{\mathrm{KL}}\big{(}{\mathbb{P}}\;\|\;{\mathbb{Q}}\big{)},$ (reverse)
Let $\|\cdot\|_{\rm{TV}}$ denote the total variation metric on path measures. We then have the following result.
Proposition B.1.
Let the iterates of our AMP be denoted by ${\mathbb{P}}^{k}$, and let the iterates of symmetrized IPFP be denoted by ${\mathbb{Q}}^{k}$. Choose a compact ball $\mathcal{C}$ in the total variation metric containing all of the iterates ${\mathbb{P}}^{k}$ and ${\mathbb{Q}}^{k}$. Suppose the following conditions hold:
  • On $\mathcal{C}$, the projection (4) is a contraction, meaning that it is Lipschitz continuous with respect to the total variation metric as a function of the second argument with Lipschitz constant $L<1$. I.e., $\|\Pi({\mathbb{P}}_{1})-\Pi({\mathbb{P}}_{2})\|_{\rm{TV}}\leq L\cdot\||{\mathbb{P}}_{1}-{\mathbb{P}}_{2}\||_{\rm{TV}}$ for every choice of path measures ${\mathbb{P}}_{1},\,{\mathbb{P}}_{2}\in\mathcal{C}$.
  • On $\mathcal{C}$, the projection (5) is non-expansive, meaning that it is Lipschitz continuous with respect to the total variation metric as a function of the second argument with Lipschitz constant $L^{\prime}\leq 1$. I.e., $\|\Phi({\mathbb{P}}_{1})-\Phi({\mathbb{P}}_{2})\|_{\rm{TV}}\leq L^{\prime}\cdot\||{\mathbb{P}}_{1}-{\mathbb{P}}_{2}\||_{\rm{TV}}$ for every choice of path measures ${\mathbb{P}}_{1},\,{\mathbb{P}}_{2}\in\mathcal{C}$.
  • For some $\varepsilon>0$, we have $\|\Phi(\Pi({\mathbb{P}}))-\Psi(\Pi({\mathbb{P}}))\|_{\rm{TV}}<\varepsilon$ for every choice of ${\mathbb{P}}\in\mathcal{C}$.
Then the total variation between the MSB and the limit of our AMP is bounded above by $\varepsilon/(1-L)$.
Proof.
We claim that the three conditions imply the following inequality:
$\|{\mathbb{P}}^{2k+3}-{\mathbb{Q}}^{2k+3}\|_{\rm{TV}}\leq L(\varepsilon+\|{\mathbb{P}}^{2k+1}-{\mathbb{Q}}^{2k+1}\|_{\rm{TV}}).$ (20)
To see why the claim holds, notice that the first condition implies
$\|{\mathbb{P}}^{2k+3}-{\mathbb{Q}}^{2k+3}\|_{\rm{TV}}\leq L(\|{\mathbb{P}}^{2k+2}-{\mathbb{Q}}^{2k+2}\|_{\rm{TV}}).$ (21)
Next, by an application of the triangle inequality followed by an application of the second and third conditions, we have
$\|{\mathbb{P}}^{2k+2}-{\mathbb{Q}}^{2k+2}\|_{\rm{TV}}$ $\leq\|\Psi({\mathbb{P}}^{2k+1})-\Phi({\mathbb{Q}}^{2k+1})\|_{\rm{TV}}$ (22)
$\leq\|\Psi({\mathbb{P}}^{2k+1})-\Phi({\mathbb{P}}^{2k+1})\|_{\rm{TV}}+\|\Phi({\mathbb{P}}^{2k+1})-\Phi({\mathbb{Q}}^{2k+1})\|_{\rm{TV}}$
$\leq\varepsilon+L^{\prime}\cdot\|{\mathbb{P}}^{2k+1}-{\mathbb{Q}}^{2k+1}\|_{\rm{TV}}.$
Combining the bounds (21) and (22) with the assumption that $L^{\prime}\leq 1$ yields (20). Solving this recursive inequality and letting $k\to\infty$ then yields the bound in the proposition. ∎
We now explain the conditions stipulated in Proposition B.1 in words. The existence of the compact ball $\mathcal{C}$ is guaranteed because both sequences of iterates converge. The first two conditions, that the projection (4) is a contraction and that the projection (5) is non-expansive, are to be expected because metric projection maps onto convex sets are non-expansive (i.e., are Lipschitz with constant $\leq 1$); note that this heuristic is not a proof of the condition, because the direct projection is computed using the $D_{\mathrm{KL}}$ divergence, which is not a metric. The Lipschitz constant $L$ depends only on the choice of the marginal $\pi$, which in our practical implementation corresponds to the data distribution.
As for the third condition, this roughly corresponds to stipulating that the reference process is not too far from the MSB, so that the reverse and direct $D_{\mathrm{KL}}$ projections of odd-numbered iterates are not too different from each other. Depending on how close $L$ is to $1$, the parameter $\varepsilon$ can be made to be sufficiently small by choosing the reference process accordingly so as to ensure that the limit of the AMP is within a prescribed distance from the MSB. If one does not have the freedom to choose the reference process but would still like to obtain an estimate of the MSB quickly, one can apply a few iterations of symmetrized IPFP to obtain an intermediary time-symmetric process $\widetilde{{\mathbb{P}}}$ that is close enough to the MSB; this intermediary process can then be used as a reference for an application of our AMP. If $\widetilde{{\mathbb{P}}}$ is sufficiently close to the MSB, then the third condition will be satisfied on the AMP iterates for a given $\varepsilon>0$. This mixed algorithm enjoys the speed improvement provided by our AMP, while retaining a prescribed level of accuracy.
Appendix C Sampling with In-Distribution Variation
Given such a sample $x_{0}\sim\pi$, we solve the SDE corresponding to the Schrödinger bridge to push $x_{0}$ forward in time, arriving at a final sample $x_{1}\in\pi$. We want $x_{1}$ to be a variation of $x_{0}$, where the proximity of $x_{1}$ to $x_{0}$ correlates with the size of the noise coefficient $\sigma$. Justifying this mathematically requires understanding how the conditional distribution ${\mathbf{X}}_{1}\mid{\mathbf{X}}_{0}=x_{0}$, specifically its mean and variance, depend on $\sigma$. While these quantities do not in general have closed form expressions, it is possible to compute them exactly in the case where $\pi={\mathcal{N}}(0,{\mathbf{I}})$ is a $1$-dimensional Gaussian.
In this case, let ${\mathbf{X}}_{t}$ be the diffusion process associated to the Schrödinger bridge, where the reference path measure corresponds to an Ornstein-Uhlenbeck process with drift coefficient $-\alpha$. In Proposition D.1 (see Appendix D for the statement and proof) we determine the joint distribution of ${\mathbf{X}}_{0}$ and ${\mathbf{X}}_{1}$ in terms of a quantity $\beta$, which is a function of $\alpha$ and $\sigma$ that grows approximately as $1+c(\alpha)\times\sigma^{2}$ for some function $c$. Let $p(x,y)$ denote the probability density function of the joint distribution of ${\mathbf{X}}_{0}$ and ${\mathbf{X}}_{1}$, and recall that $p(x,y)$ is the product of the conditional PDF of ${\mathbf{X}}_{1}\mid{\mathbf{X}}_{0}$ with the PDF of ${\mathbf{X}}_{0}$. Using this fact in conjunction with Proposition D.1, the PDF of ${\mathbf{X}}_{1}\mid{\mathbf{X}}_{0}=x_{0}$ is
$\frac{p(x_{0},y)}{p_{{\mathbf{X}}_{0}}(x_{0})}=\textrm{exp}\bigg{(}{-\frac{1}{2(1-\beta^{2})}(x_{0}^{2}-2\beta x_{0}y+y^{2})+\frac{x_{0}^{2}}{2}}\bigg{)}.$
From the right-hand side, we see that ${\mathbf{X}}_{1}\mid{\mathbf{X}}_{0}=x_{0}$ is Gaussian with mean $x_{0}(\nicefrac{{\beta}}{{1-\beta^{2}}})$ and variance $1-\beta^{2}$. Thus, changing the noise value $\sigma$ alters both the mean and variance of samples pushed forward via the Schrödinger bridge. Indeed, in the case of the mean, it grows inversely proportional to $\sigma^{2}$. Consequently, if $\sigma<1$, then we should expect the Schrödinger bridge to push samples away from the distribution mean, whereas if $\sigma>1$, then the opposite occurs, and samples experience mean reversion. As for the variance, note that $1-\beta^{2}$ grows at least as fast as $\sigma^{2}$, so we should expect the Schrödinger bridge to produce samples with spread that increases as $\sigma$ increases. We expect that similar effects occur even when the marginal distribution $\pi$ is not Gaussian: i.e., the value of $\sigma$ should be directly related to the proximity of generated samples in an analogous way. As our method gives an approximation of the true mirror Schrödinger bridge, we expect the mirror bridge to inherit these properties.
Appendix D Analytical Solution to MSB in the Gaussian Case
Proposition D.1.